# Quantum Bit Threads and the Entropohedron

### v1

@@ 1 Introduction
Holographic entanglement entropy, a cornerstone of our modern understanding of quantum gravity, reveals how bulk geometry encodes boundary entropy, and vice versa. There are two equivalent sets of prescriptions for holographic entanglement entropy, the older being surface-based, starting with the Ryu-Takayanagi (RT) formula Ryu:2006bv .
@@
The more recent formulation of holographic entanglement entropy are the bit thread prescriptions. A classical bit thread is a continuous bulk curve connecting $A$ to its complement $A^{c}$. In the bit thread reformulation of the RT formula Freedman:2016zud , $S(A)$ equals the maximum number of threads connecting $A$ to $A^{c}$, subject to a bound on their density, $\rho\leq 1/4G_{\rm N}$. Classical bit threads can be described mathematically as the field lines of a divergenceless bulk vector field $v$. With the density of threads identified with the norm of $v$, the density bound becomes a norm bound $|v|\leq 1/4G_{\rm N}$ and the number of threads connecting $A$ to $A^{c}$ is identified with the flux of $v$ on $A$, giving the classical flow formula:
@@
where $n$ is the inward-directed unit normal vector. The equivalence between (2) and the original RT formula (1) is guaranteed by the Riemannian max flow-min cut theorem federer1974real ; strang1983maximal ; sullivan1990crystalline ; nozawa1990max ; Headrick:2017ucz . With bit threads, the connection between boundary entanglement and bulk geometry is even more manifest than in the RT formula. Bit thread prescriptions have both conceptual advantages over surface-based prescriptions, such as making the physical meaning of entropy inequalities manifest, and technical advantages, as they are expressed as convex optimisation problems.
@@
The QES formula has led to important insights in quantum gravity, such as, in the context of black hole evaporation, entropy curves consistent with unitarity, and the presence of entanglement islands Almheiri:2019psf ; Penington:2019npb ; Almheiri_2021 .
There are a number of subtle and interesting issues related to both the definition of the generalized entropy (such as the contribution of the gravitational field itself Colin-Ellerin:2025dgq and additional terms that depend on the details and dimensionality of the bulk theory), and the QES formula (such as conditions on the state of the bulk fields for its validity Akers:2020pmf ). Nonetheless, in this paper, we will focus on the QES formula in the form (3), as it captures the essence of the bulk entropy contribution and is rich enough to be worth studying on its own.
A bit thread reformulation of the QES formula (3) was given in the paper Rolph:2021hgz . [^2] Other bit-thread formulas including quantum corrections to $S(A)$ have been proposed. For example, the paper Agon:2021tia gave a flow proposal that captures corrections to first order in $G_{\rm N}$ (whereas the QES formula and Rolph:2021hgz account for correction to all order in $G_{\rm N}$). And Du:2024xoz gave a flow reformulation of the “generalised entanglement wedge” proposal of Bousso:2022hlz . Specifically, the definition of a flow was relaxed to allow for a non-zero divergence controlled by the bulk entropy:
@@
For reasons that we will shortly make clear, we call this the loose quantum flow prescription. In this quantum prescription, unlike the classical case, bit threads can start and end at points in the bulk, with the number of threads that can end in a given bulk region bounded by the region’s entropy. This allows for more flux to pass through $A$ than can pass through any surface homologous to $A$. The paper Rolph:2021hgz proved the equivalence of (5) and (3), and explored its properties in general as well as in several examples. Note that the data required to evaluate (5) is precisely the same as that required to evaluate (3), namely the metric and the entropy of every bulk region coincident to $A$. It’s also worth noting that the proof of equivalence relied essentially on the strong subadditivity property of the bulk entropies.
@@
In this paper, we will extend the work of Rolph:2021hgz with more bit thread reformulations of the QES formula (3). First, we will show that it is possible to impose a much stronger constraint on the divergence of $v$ without changing the maximum flux, giving the strict quantum flow prescription:
@@
The constraint in (6) on the divergence of $v$ differs from that in (5) in two ways: first, it is applied to every bulk region $r$, with $\mathcal{R}$ defined as the set of all bulk regions, without any condition on their intersection with the boundary; second, the integrated divergence is bounded above by $S_{\rm b}(r)$ as well as below by $S_{\rm b}(r)$. Both constraints are non-local and bound the flux into bulk subregions; the larger the bulk entropy of the region, the larger the number of bit threads that can end there. Unlike the loose constraint, but like classical flows, the strict constraint is independent of the choice of boundary region $A$. We will show that it is automatically obeyed in doubly holographic systems; for this purpose, we will prove a new theorem giving necessary and sufficient conditions for a function on the boundary of a Riemannian manifold to be the boundary flux of a classical flow. More generally, the strict divergence constraint leads to an interesting generalization of the notion of an entanglement contour Vidal:2014aal , which we call an entanglement distribution function (EDF). For a finite set of $N$ parties, the set of all EDFs for a given state is a convex polyhedron in $\mathbf{R}^{N}$ which we call the entropohedron. We will determine its properties and study how it reflects the entanglement structure of the state. We also consider the generalization of strict quantum flows to multiflows.
From the bit thread perspective, entanglement islands are highly entangled bulk regions where so many threads are forced to reappear because of the divergence constraint that the threads become maximally packed on a surface disconnected from the asymptotic boundary, and this bottleneck is the boundary of the island; see Fig. 1. In Fig. 2, we depict an example where the AdS bulk is entangled with a closed universe, and the quantum bit threads can jump across, but they have to return because the constraint in (6) requires that $\int_{\Sigma}\nabla\cdot v=0$ for a pure boundary state. For a given setup, there are generically many optimal flow configurations, and $v=0$ is the simplest for the example shown in Fig. 2. In Fig. 3, inspired by Antonini:2023hdh , we depict an example where optimal quantum thread configurations in (6) must pass through an entangled closed universe on their way between the two AdS boundaries.
@@
As the flow field is divergenceless, this prescription is, in a sense, the closest to the original classical formulation of Freedman:2016zud .
An alternative mathematical formulation of bit threads, instead of flows, is in terms of so-called thread distributions Headrick:2022nbe . In the classical case, a thread is a continuous bulk curve connecting boundary points, and a thread distribution is a measure on the set of threads subject to a density bound. Inspired by the quantum flows, we propose a quantum thread distribution formula, in which the quantum threads are allowed to jump from one point to another of the bulk, subject to the constraint that the number of threads leaving or entering any bulk region is bounded above by its entropy. We give some evidence, in particular based on double holography, that this thread distribution formula is equivalent to the QES formula. However, several statements concerning quantum bit thread distributions are left as conjectures.
In the form of the QES prescription given in (3), we are accounting for bulk quantum corrections but, implicitly, staying in a static or time reflection-symmetric setting — as we will throughout this paper — where the extremization part of the QES formula simplifies to a minimization over subregions of a single time slice. In a forthcoming paper CQBT , we will drop the static assumption, and derive covariant quantum bit thread prescriptions, which synthesize and extend the quantum bit threads with the covariant classical bit threads of Headrick:2022nbe , which are equivalent to the covariant Hubeny-Rangamani-Takayanagi (HRT) formula.
@@ 2.1 Flows in double holography
A doubly holographic theory is one where the bulk matter fields are themselves holographic, described by a higher-dimensional second bulk Karch:2000gx ; Karch:2000ct ; Emparan:2002px ; Almheiri:2019hni . Typically, it is assumed that the theory on the second bulk may be treated classically, hence that the RT formula applies for computing entropies of regions in the first bulk. The geometric nature of these bulk entropies makes the computation of entropies in doubly holographic theories more tractable than in generic holographic theories, and is a source of useful intuition. We will use double holography to motivate the definition of strict quantum flows.
@@ 2.2 Definition & basic properties of strict flows
We now consider a general holographic theory, in which $S(A)$ is computed by the QES formula (3), with the bulk entropy $S_{\rm b}$ being a function on the set $\mathcal{R}$ of bulk regions that obeys positivity and strong subadditivity but is otherwise general. Motivated by what we found in doubly holographic theories in the previous subsection, we define a strict quantum flow as a vector field $v$ on $\Sigma$ obeying (32). This definition should be compared with that of the quantum flows defined in Rolph:2021hgz , which we will call “loose”. [^3] One can also consider intermediate definitions, with $\mathcal{R}_{A}$ and the absolute value, or with $\mathcal{R}$ and no absolute value. Since loose and strict flows both have maximum flux equal to $S(A)$, these intermediate flows must also have this property. A loose quantum flow for a boundary region $A$ is a vector field $v$ on $\Sigma$ such that
@@
where $\mathcal{R}_{A}$ is the set of regions $r\subseteq\Sigma$ such that $r\cap\partial\Sigma=A$. The differences are, first, the divergence bound is applied only to regions in $\mathcal{R}_{A}$, and second, the integrated divergence is bounded only in one direction. In Rolph:2021hgz , it was shown that the maximum flux through $A$ for a loose flow equals $S(A)$. The same holds for strict flows:
@@
See theorem 2.4 for the proof of (34). In fact, we will give two proofs, both using Lagrange dualization: one dualizing from (34) to the QES prescription, and one doing the reverse. Flow-to-surface dualizations are useful when one has a bit thread proposal in hand, like (34), that one wishes to prove is equivalent to the corresponding surface-based prescription, while surface-to-flow dualizations are useful when, such as for the covariant QES prescription, there is not an obvious bit thread proposal, and one wishes to derive it starting from the surface-based prescription. So, the pair of proofs illustrate how the dualization works in both directions, and the surface-to-flow dualization will be useful for the covariant setup CQBT . Also, in theorem 2.5, we will also give a new proof that the maximum flux over loose quantum flows equals $S(A)$. This proof is complementary to the proof given in Rolph:2021hgz , because the Lagrange dualization is in the reverse direction, and it is complementary to our second proof of theorem 2.4 because it smears out bulk regions a different way, using level sets rather than measures. Notably, all of these proofs, as well as those of theorems 2.6 and 2.7, make use of the assumption that $S_{\rm b}$ obeys strong subadditivity (SSA).
@@
Because of the flow’s divergence constraint, in particular the absolute value, the strict flow satisfies $\int_{\Sigma}\nabla\cdot v=0$, assuming that the bulk state is pure, which means the net number of flow lines that end in any bulk region $r$ equals the net number that start in $r^{c}$. This is important for the pictorial interpretation that the quantum bit threads jump across the QES. In contrast, with the loose divergence constraint, max-flow lines end in the entanglement wedge of $A$ but do not have to reappear in the complement region.
@@ 2.3 Entanglement distribution functions
We define an entanglement distribution function (EDF) [^4] EDFs are closely related to the notions of entanglement contour and partial entanglement entropy Vidal:2014aal ; Wen:2019iyq ; Kudler-Flam:2019oru ; Rolph:2021nan . Given a region $A$, these are non-negative function $s_{A}$ on $A$ obeying $s_{A}\geq 0$, $\int_{A}s_{A}=S(A)$, and several other conditions. In our case, since we are imposing (35) for all regions, we must allow $f$ to be negative; otherwise, for example, if the full state is pure, the only solution would be $f=0$. However, as we will see, for any fixed region $A$, an EDF can be found that is non-negative on $A$ and integrates to $S(A)$. For a discussion of entanglement contours, see Rolph:2021nan . Other related notions include various quantities that have been given the name “entanglement density” Nozaki:2013wia ; Bhattacharya:2014vja ; Gushterov:2017vnr ; Jeong:2022zea . as a function $f$ on $M$ such that, for all $A\in\mathcal{A}$,
@@ 2.4 More on strict flows
For the proof, see theorem 2.7. (Nesting was proven for loose flows in Rolph:2021hgz .)
Alternatively, nesting follows as a corollary of two theorems quoted above: an EDF exists that saturates on any nested set of regions (see subsection 2.3); and a strict flow exists whose boundary flux equals any given EDF (see just above). However, the latter route assumes that $S$ obeys SSA (used in proving the existence of a saturating EDF), whereas the direct proof of the theorem on nesting of strict flows does not. The distinction is noteworthy because nesting can itself be used to prove SSA, by the following argument Freedman:2016zud . Let $A,B,C$ be disjoint regions, and let $v$ be a max flow for both $B$ and $ABC$. Then we have $S(AB)\geq\int_{AB}n\cdot v$, $S(BC)\geq\int_{BC}n\cdot v$, so
@@
Classical flows also obey a stronger property, the existence of a so-called max multiflow Cui:2018dyq , which implies that $S$ obeys the monogamy of mutual information (MMI) inequality Hayden:2011ag ,

### v2

@@ 1 Introduction
Holographic entanglement entropy, a cornerstone of our modern understanding of quantum gravity, reveals how bulk geometry encodes boundary entropy, and vice versa. There are two equivalent sets of prescriptions for holographic entanglement entropy, the older being surface-based, starting with the Ryu-Takayanagi (RT) formula Ryu and Takayanagi (2006).
@@
The more recent formulation of holographic entanglement entropy are the bit thread prescriptions. A classical bit thread is a continuous bulk curve connecting $A$ to its complement $A^{c}$. In the bit thread reformulation of the RT formula Freedman and Headrick (2017), $S(A)$ equals the maximum number of threads connecting $A$ to $A^{c}$, subject to a bound on their density, $\rho\leq 1/4G_{\rm N}$. Classical bit threads can be described mathematically as the field lines of a divergenceless bulk vector field $v$. With the density of threads identified with the norm of $v$, the density bound becomes a norm bound $|v|\leq 1/4G_{\rm N}$ and the number of threads connecting $A$ to $A^{c}$ is identified with the flux of $v$ on $A$, giving the classical flow formula:
@@
where $n$ is the inward-directed unit normal vector. The equivalence between (2) and the original RT formula (1) is guaranteed by the Riemannian max flow-min cut theorem Federer (1974); Strang (1983); Sullivan (1990); Nozawa (1990); Headrick and Hubeny (2018). With bit threads, the connection between boundary entanglement and bulk geometry is even more manifest than in the RT formula. Bit thread prescriptions have both conceptual advantages over surface-based prescriptions, such as making the physical meaning of entropy inequalities manifest, and technical advantages, as they are expressed as convex optimisation problems.
@@
The QES formula has led to important insights in quantum gravity, such as, in the context of black hole evaporation, entropy curves consistent with unitarity, and the presence of entanglement islands Almheiri et al. (2019); Penington (2020); Almheiri et al. (2021).
There are a number of subtle and interesting issues related to both the definition of the generalized entropy (such as the contribution of the gravitational field itself Colin-Ellerin et al. (2025) and additional terms that depend on the details and dimensionality of the bulk theory), and the QES formula (such as conditions on the state of the bulk fields for its validity Akers and Penington (2021)). Nonetheless, in this paper, we will focus on the QES formula in the form (3), as it captures the essence of the bulk entropy contribution and is rich enough to be worth studying on its own.
A bit thread reformulation of the QES formula (3) was given in the paper Rolph (2023). [^2] Other bit-thread formulas including quantum corrections to $S(A)$ have been proposed. For example, the paper Agón and Pedraza (2022) gave a flow proposal that captures corrections to first order in $G_{\rm N}$ (whereas the QES formula and Rolph (2023) account for correction to all order in $G_{\rm N}$). And Du and Sun (2024) gave a flow reformulation of the “generalised entanglement wedge” proposal of Bousso and Penington (2023). Specifically, the definition of a flow was relaxed to allow for a non-zero divergence controlled by the bulk entropy:
@@
For reasons that we will shortly make clear, we call this the loose quantum flow prescription. In this quantum prescription, unlike the classical case, bit threads can start and end at points in the bulk, with the number of threads that can end in a given bulk region bounded by the region’s entropy. This allows for more flux to pass through $A$ than can pass through any surface homologous to $A$. The paper Rolph (2023) proved the equivalence of (5) and (3), and explored its properties in general as well as in several examples. Note that the data required to evaluate (5) is precisely the same as that required to evaluate (3), namely the metric and the entropy of every bulk region coincident to $A$. It’s also worth noting that the proof of equivalence relied essentially on the strong subadditivity property of the bulk entropies.
@@
In this paper, we will extend the work of Rolph (2023) with more bit thread reformulations of the QES formula (3). First, we will show that it is possible to impose a much stronger constraint on the divergence of $v$ without changing the maximum flux, giving the strict quantum flow prescription:
@@
The constraint in (6) on the divergence of $v$ differs from that in (5) in two ways: first, it is applied to every bulk region $r$, with $\mathcal{R}$ defined as the set of all bulk regions, without any condition on their intersection with the boundary; second, the integrated divergence is bounded above by $S_{\rm b}(r)$ as well as below by $S_{\rm b}(r)$. Both constraints are non-local and bound the flux into bulk subregions; the larger the bulk entropy of the region, the larger the number of bit threads that can end there. Unlike the loose constraint, but like classical flows, the strict constraint is independent of the choice of boundary region $A$. We will show that it is automatically obeyed in doubly holographic systems; for this purpose, we will prove a new theorem giving necessary and sufficient conditions for a function on the boundary of a Riemannian manifold to be the boundary flux of a classical flow. More generally, the strict divergence constraint leads to an interesting generalization of the notion of an entanglement contour Vidal and Chen (2014), which we call an entanglement distribution function (EDF). For a finite set of $N$ parties, the set of all EDFs for a given state is a convex polyhedron in $\mathbf{R}^{N}$ which we call the entropohedron. We will determine its properties and study how it reflects the entanglement structure of the state. We also consider the generalization of strict quantum flows to multiflows.
From the bit thread perspective, entanglement islands are highly entangled bulk regions where so many threads are forced to reappear because of the divergence constraint that the threads become maximally packed on a surface disconnected from the asymptotic boundary, and this bottleneck is the boundary of the island; see Fig. 1. In Fig. 2, we depict an example where the AdS bulk is entangled with a closed universe, and the quantum bit threads can jump across, but they have to return because the constraint in (6) requires that $\int_{\Sigma}\nabla\cdot v=0$ for a pure boundary state. For a given setup, there are generically many optimal flow configurations, and $v=0$ is the simplest for the example shown in Fig. 2. In Fig. 3, inspired by Antonini et al. (2023), we depict an example where optimal quantum thread configurations in (6) must pass through an entangled closed universe on their way between the two AdS boundaries.
@@
As the flow field is divergenceless, this prescription is, in a sense, the closest to the original classical formulation of Freedman and Headrick (2017).
An alternative mathematical formulation of bit threads, instead of flows, is in terms of so-called thread distributions Headrick and Hubeny (2023). In the classical case, a thread is a continuous bulk curve connecting boundary points, and a thread distribution is a measure on the set of threads subject to a density bound. Inspired by the quantum flows, we propose a quantum thread distribution formula, in which the quantum threads are allowed to jump from one point to another of the bulk, subject to the constraint that the number of threads leaving or entering any bulk region is bounded above by its entropy. We give some evidence, in particular based on double holography, that this thread distribution formula is equivalent to the QES formula. However, several statements concerning quantum bit thread distributions are left as conjectures.
In the form of the QES prescription given in (3), we are accounting for bulk quantum corrections but, implicitly, staying in a static or time reflection-symmetric setting — as we will throughout this paper — where the extremization part of the QES formula simplifies to a minimization over subregions of a single time slice. In a forthcoming paper Headrick et al. , we will drop the static assumption, and derive covariant quantum bit thread prescriptions, which synthesize and extend the quantum bit threads with the covariant classical bit threads of Headrick and Hubeny (2023), which are equivalent to the covariant Hubeny-Rangamani-Takayanagi (HRT) formula.
@@ 2.1 Flows in double holography
A doubly holographic theory is one where the bulk matter fields are themselves holographic, described by a higher-dimensional second bulk Karch and Randall (2001b, a); Emparan et al. (2002); Almheiri et al. (2020). Typically, it is assumed that the theory on the second bulk may be treated classically, hence that the RT formula applies for computing entropies of regions in the first bulk. The geometric nature of these bulk entropies makes the computation of entropies in doubly holographic theories more tractable than in generic holographic theories, and is a source of useful intuition. We will use double holography to motivate the definition of strict quantum flows.
@@ 2.2 Definition & basic properties of strict flows
We now consider a general holographic theory, in which $S(A)$ is computed by the QES formula (3), with the bulk entropy $S_{\rm b}$ being a function on the set $\mathcal{R}$ of bulk regions that obeys positivity and strong subadditivity but is otherwise general. Motivated by what we found in doubly holographic theories in the previous subsection, we define a strict quantum flow as a vector field $v$ on $\Sigma$ obeying (32). This definition should be compared with that of the quantum flows defined in Rolph (2023), which we will call “loose”. [^3] One can also consider intermediate definitions, with $\mathcal{R}_{A}$ and the absolute value, or with $\mathcal{R}$ and no absolute value. Since loose and strict flows both have maximum flux equal to $S(A)$, these intermediate flows must also have this property. A loose quantum flow for a boundary region $A$ is a vector field $v$ on $\Sigma$ such that
@@
where $\mathcal{R}_{A}$ is the set of regions $r\subseteq\Sigma$ such that $r\cap\partial\Sigma=A$. The differences are, first, the divergence bound is applied only to regions in $\mathcal{R}_{A}$, and second, the integrated divergence is bounded only in one direction. In Rolph (2023), it was shown that the maximum flux through $A$ for a loose flow equals $S(A)$. The same holds for strict flows:
@@
See theorem 2.4 for the proof of (34). In fact, we will give two proofs, both using Lagrange dualization: one dualizing from (34) to the QES prescription, and one doing the reverse. Flow-to-surface dualizations are useful when one has a bit thread proposal in hand, like (34), that one wishes to prove is equivalent to the corresponding surface-based prescription, while surface-to-flow dualizations are useful when, such as for the covariant QES prescription, there is not an obvious bit thread proposal, and one wishes to derive it starting from the surface-based prescription. So, the pair of proofs illustrate how the dualization works in both directions, and the surface-to-flow dualization will be useful for the covariant setup Headrick et al. . Also, in theorem 2.5, we will also give a new proof that the maximum flux over loose quantum flows equals $S(A)$. This proof is complementary to the proof given in Rolph (2023), because the Lagrange dualization is in the reverse direction, and it is complementary to our second proof of theorem 2.4 because it smears out bulk regions a different way, using level sets rather than measures. Notably, all of these proofs, as well as those of theorems 2.6 and 2.7, make use of the assumption that $S_{\rm b}$ obeys strong subadditivity (SSA).
@@
Because of the flow’s divergence constraint, in particular the absolute value, the strict flow satisfies $\int_{\Sigma}\nabla\cdot v=0$, assuming that the bulk state is pure, which means the net number of flow lines that end in any bulk region $r$ equals the net number that start in $r^{c}$. This is important for the pictorial interpretation that the quantum bit threads jump across the QES. In contrast, with the loose divergence constraint, max-flow lines end in the entanglement wedge of AA [^4] Throughout this paper, what we mean by the entanglement wedge of $A$ is the intersection of the time-reflection-symmetric bulk time slice with the common definition of the entanglement wedge of $A$ (the domain of dependence of any codim-1 spacelike surface whose boundary is $A\cup\text{QES}_{A}$). Our entanglement wedge is a codim-0 region with respect to the bulk time slice. but do not have to reappear in the complement region.
@@ 2.3 Entanglement distribution functions
We define an entanglement distribution function (EDF) [^5] EDFs are closely related to the notions of entanglement contour and partial entanglement entropy Vidal and Chen (2014); Wen (2020); Kudler-Flam et al. (2019); Rolph (2022). Given a region $A$, these are non-negative function $s_{A}$ on $A$ obeying $s_{A}\geq 0$, $\int_{A}s_{A}=S(A)$, and several other conditions. In our case, since we are imposing (35) for all regions, we must allow $f$ to be negative; otherwise, for example, if the full state is pure, the only solution would be $f=0$. However, as we will see, for any fixed region $A$, an EDF can be found that is non-negative on $A$ and integrates to $S(A)$. For a discussion of entanglement contours, see Rolph (2022). Other related notions include various quantities that have been given the name “entanglement density” Nozaki et al. (2013); Bhattacharya et al. (2015); Gushterov et al. (2017); Jeong et al. (2022). as a function $f$ on $M$ such that, for all $A\in\mathcal{A}$,
@@ 2.4 More on strict flows
For the proof, see theorem 2.7. (Nesting was proven for loose flows in Rolph (2023).)
Alternatively, nesting follows as a corollary of two theorems quoted above: an EDF exists that saturates on any nested set of regions (see subsection 2.3); and a strict flow exists whose boundary flux equals any given EDF (see just above). However, the latter route assumes that $S$ obeys SSA (used in proving the existence of a saturating EDF), whereas the direct proof of the theorem on nesting of strict flows does not. The distinction is noteworthy because nesting can itself be used to prove SSA, by the following argument Freedman and Headrick (2017). Let $A,B,C$ be disjoint regions, and let $v$ be a max flow for both $B$ and $ABC$. Then we have $S(AB)\geq\int_{AB}n\cdot v$, $S(BC)\geq\int_{BC}n\cdot v$, so
@@
Classical flows also obey a stronger property, the existence of a so-called max multiflow Cui et al. (2019), which implies that $S$ obeys the monogamy of mutual information (MMI) inequality Hayden et al. (2013),